13. Test at the 0.01 level the statement that 55% of those families who plan to purchase a vacation residence in Florida want a condominium. The null hypothesis is π = .55 and the alternate is π ≠ .55. A random sample of 300 families who planned to buy a vacation residence revealed that 218 families want a condominium. What decision should be made regarding the null hypothesis?
A. Reject it B. Do not reject it C. Cannot accept nor reject it based on the information given D. None of the above

Answer by Boreal(15235)

Ho:p=0.55
Ha: pNE 0.55
alpha=0.05 p{reject Ho|Ho true}=0.01
test is a 1-sample proportion 2 sided, with critical value being |z|>2.576
p hat=218/300=0.727
z=(phat-p)/sqrt(0.55*0.45/300)
=0.177/0.0287
6.16
reject Ho, choice A.